what is the oxidation state of iodine in iodate ion?
1
Expert's answer
2015-09-15T05:05:39-0400
Iodate ion: empirical formulae is IO3-
The oxidation state of Oxygen is equal to -2. Then with easy calculation we get the oxidation state of iodine:
Ox(I) + Ox(O)*3 = -1
OX(I) = -1 - (Ox(O)*3) = -1 +2*3 = 5